OpenStudy (jazzyluvsu123):

is 78.111 / 3= 26.033333333...... @agent0smith @Agent47 @mathtutoring22 @UnkleRhaukus @sourwing

OpenStudy (agent0smith):

check with a calculator..? Computers have one. Google too.

OpenStudy (jazzyluvsu123):

okk

OpenStudy (unklerhaukus):

78.111 / 3 = (75+3+0.111)/3 = 75/3 +3/3+0.111/3 =

OpenStudy (jazzyluvsu123):

= ? @UnkleRhaukus

OpenStudy (unklerhaukus):

= 25+1+0.111/3

OpenStudy (jazzyluvsu123):

= 26.037 right ?

OpenStudy (unklerhaukus):

yes